OpenStudy (anonymous):

1/6-3/2x=1/5x solve algebraically Help please

OpenStudy (anonymous):

kuch bhi ata nhi kya

OpenStudy (anonymous):

What?

OpenStudy (anonymous):

easiest qn

OpenStudy (kropot72):

Is the question: \[1/6-3/2(x)=1/5(x)\]

OpenStudy (anonymous):

yes

OpenStudy (kropot72):

Multiplying both sides by 10 gives: 10/6 - 15x = 2x 10/6 = 17x x = 10/(6 * 17) = 10/102 = 5/51

OpenStudy (precal):